When the accounts of Vaughn Inc. are examined, the adjusting data listed below are uncovered on December 31, the end of an annual fiscal period.

1. The prepaid insurance account shows a debit of $4,800, representing the cost of a 2-year fire insurance policy dated August 1 of the current year.
2. On November 1, Rent Revenue was credited for $1,869, representing revenue from a subrental for a 3-month period beginning on that date.
3. Purchase of advertising materials for $845 during the year was recorded in the Advertising Expense account. On December 31, advertising materials of $295 are on hand.
4. Interest of $721 has accrued on notes payable. The interest will be paid in January of the next year.


A.) Prepare the following in general journal form.


(b) The reversing entry for each item where appropriate.
Answer #1

All amounts are in $

A. Journal Entries

1.

Insurance expense $1,000

Prepaid Insurance $1,000

(Insurance expense for 5 months from August to December recorded)

Insurance expense = (4,800/24) x 5 = 1,000

2.

Rent Revenue $289.67

Unearned Rent revenue $289.67

(Rent Revenue for January month transferred to unearned Rent revenue)

Unearned Rent Revenue = 869/3 = 289.67

3.

Advertising materials $295

Advertising expenses $295

(Unused advertising materials accounted by reducing the advertising expense)

4.

Interest Expense $721

Accrued interest $721

(Accrued interest accounted)

B. Reversing journal entries

1.

No reversing entry is required, as we accounted the pre-paid insurance using "Asset method" and not "Expense method"

2.

Unearned Rent Revenue $289.67

Rent Revenue $289.67

(Reversing entry for unearned revenue accounted - because "income method" is used to account and not "liability method")

3.

No reversing journal entry is required. Expense entry is recorded in the period of actual usage

4.

Accrued interest $721

Interest expense $721

(Reversing entry for interest expense)

Note :

Instead of accrued interest account, Interest payable account can also be used